OpenStudy (anonymous):

A body projected Vertically upwards with a velocity v at t=0 is found at a height h after 1 second and after a further 6 second it is found at the same height (g=10m/s^2) then h=30m u=40 Hmax=80m distance moved in 5th second is 5m this is a multiple answer question more than one answer is correct.

OpenStudy (ajprincess):

vat is u here?

OpenStudy (anonymous):

Not given?

OpenStudy (ajprincess):

i thnk it must b v.

OpenStudy (anonymous):

all are correct

OpenStudy (anonymous):

@Arnab09 no answer is b c d

OpenStudy (anonymous):

okay, lemme check again

OpenStudy (anonymous):

if b is true, a is ought to be true

OpenStudy (anonymous):

No......@Arnab09

OpenStudy (anonymous):

lemme check again, h=ut- 1/2 gt^2.. right?

OpenStudy (anonymous):

what a fool i am.. h has to be 35 m

OpenStudy (anonymous):

b,c,d why r u posting these questions, @open_study1 ?

OpenStudy (anonymous):

I wanted u knw the steps to do this

OpenStudy (anonymous):

apply basic formulae, thats all

OpenStudy (anonymous):

only trick is here: after 6 seconds it returns to the same height. so, it will take 6/2=3 seconds more to reach the Hmax.. so, it will take 4 seconds to reach Hmax at Hmax, v=0 so, u can apply 0=u-gt find u and u can do the rest, i believe

OpenStudy (anonymous):

@Arnab09 i will try my best to solve this plzz wait and will inform u..

OpenStudy (anonymous):

i got Hmax=80m

OpenStudy (anonymous):

and u as 40m

OpenStudy (anonymous):

Distance in 5 second ???

OpenStudy (anonymous):

5th second

OpenStudy (anonymous):

u+ 1/2 a(2t-1) u will find it by calculation

OpenStudy (anonymous):